OpenStudy (anonymous):

a 12 foot ladder is resting against a wall. the base of the ladder is 2.5 feet from the base of the wall. how high up the wall will the ladder reach

OpenStudy (anonymous):

a^2 + b^2 = c ^2 2.5^2 + b^2 = 12^2 6.25 + b^2 = 144 minus 6.25 b^2 = 137.75 take the square root b = 11.736694543 feet if you need to round 12 or 11.7

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.